# A Disproof of Large Language Model Consciousness: The Necessity of Continual Learning for Consciousness

### v1

@@ 4 The Proximity Argument
While there exist in the literature previous “no-go” results on AI consciousness, such as arguments against AI consciousness based on causal exclusion [kleiner2024dynamically] (though similar arguments have been given against the efficacy of human consciousness as well [kim2000mind]), as well as other arguments for biological naturalism [seth2024conscious, campero2025consciousness], one unique aspect of the Proximity Argument is that it makes minimal assumptions other than that: (a) a theory of consciousness cannot be trivial or a priori falsified, (b) that certain universal substitutions are possible (which we know from existent mathematical theorems to be true), and (c) that failed predictions (in the form of widespread or universal mismatches) from a theory in the comparative testing setup described would falsify that theory. Nor does it rely on a particular theory; e.g., while transformers provably are entirely feedforward and have zero integrated information (thus, IIT would say that LLMs are not conscious) [ali2025intelligence], the Proximity Argument goes beyond relying on a particular theory.
E.g., the Proximity Argument requires no opinion on whether forms of “functionalism” are true or false, or if qualia is causally relevant or epiphenomenal, nor any other philosophical position about consciousness, nor does it rely on any particular theory of consciousness or even class of theories to be true (it does not, e.g., require solving the Hard Problem [chalmers1995facing]. It is a philosophically neutral proof that both constrains and disproves LLM consciousness based on falsification. All the Proximity Argument requires is that the Kleiner-Hoel dilemma apply in the transitive chain of substitutions from an LLM to systems for which there are no non-trivial theories of consciousness available. There are very good reasons to believe that it does. E.g., the links of this transitive chain are proximal in “substitution space” since we know via the universal approximation theorem [hornik1989multilayer] that various artificial neural networks of much the same substrate and workings to LLMs are (a) either non-conscious themselves, or (b) are in turn proximal to non-conscious systems like lookup tables.
At absolute minimum, the Proximity Argument means that the space for theories of consciousness to grant consciousness to LLMs is highly constrained (Theorem 4.6) to what is lost in this definable chain of substitutions (which does not seem appropriate to ground a rich and non-trivial theory of consciousness in). However, given the further Proposition 4.8, we can say that LLMs are non-conscious, since the very act of there being a well-defined substitution chain means that for any non-trivial theory such substitutions would also necessarily create mismatches and so lead to a priori falsification. A specific example of how this works, and why compression cannot be a ground of a theory of consciousness, is given in the Corollary 4.10 proof that single-hidden-layer FNNs are themselves non-conscious systems.
4.1 Why the Proximity Argument applies in LLMs but not in humans
@@
The Proximity Argument is based on the formal falsification framework developed in [kleiner2021falsification]; however, the Proximity Argument does not require the full acceptance of the Kleiner-Hoel dilemma applying in humans or for all possible theories of consciousness. Humans are much less proximate (Proposition 4.5) in “substitution distance” (Definition 4.3) to trivially non-conscious systems, therefore the Constraint Theorem (Theorem 4.6) doesn’t bound the possibility space to the same degree, as there are many more properties lost in universal substitutions. This means there are far more available choices to ground non-trivial theories of consciousness in.
Additionally, in the substitution chain of the Proximity Argument against LLM consciousness, one potential (if still debatable) objection to the Kleiner-Hoel dilemma doesn’t apply: That when it comes to testing theories of consciousness in humans we might weight our inferences differently for humans vs. other systems [albantakis2020unfolding]. E.g., given a universal substitution for a human in an experimental setup, we might simply decide not to falsify a theory of consciousness following universal mismatches, if we simply don’t trust the reports or behavior of that system to contain any (or barely any) scientific information about consciousness. Note that this asymmetry vanishes entirely in the Proximity Argument. When it comes to LLMs vs. trivially non-conscious systems acting as universal substitutions for them, all (or the vast majority) of the force of our inferences about consciousness comes from their complex and intelligent inputs/outputs. Yet the same thing exists for trivially non-conscious systems like lookup tables, since they necessarily share those inputs/outputs.
Finally, and by far (potentially) most importantly, there may be examples of lenient dependency that apply in humans but not LLMs. If so, then the interpretation at the heart of the definition 4.2 of Non-conscious Systems (that if a system is constrained to trivial theories, it must be non-conscious) becomes certain.
@@
Therefore, the second part of this paper will explore the requirements for lenient dependency in humans. First, I argue that lenient dependency must be a continually-existing criterion for theories of consciousness. Then, I propose that learning is a form of lenient dependency in humans, and that this is necessarily restricted to truly continual learning (the learning occurring with, or potentially even constituting, every single conscious experience). If this further argument is correct, it would significantly constrain consciousness in humans to be about the process of continual learning. However, I argue that this restriction is scientifically positive: while it obviates most existent theories in the field, it still points us toward evocative and rich theories of consciousness. Additionally, it shows that a fundamental difference between LLMs and humans is the ability to continually learn and that LLM’s lack of consciousness may be tied to this inability.
5 The Necessity of Continual Lenient Dependency for Theories of Consciousness
Before investigating a specific proposed case of lenient dependency in the form of learning, there is a broader question: given a case of lenient dependency based on some properties, how much does this constrain our current theories of consciousness?
@@ 5 The Necessity of Continual Lenient Dependency for Theories of Consciousness
Here, I argue that it forms a strong constraint on theories, in that theories must constantly be based on properties that satisfy lenient dependency at every prediction by the theory with regard to the contents or presence of consciousness. In other words, lenient dependency is a continual and ongoing requirement for theories of consciousness.
@@
Let $T$ be a theory of consciousness with lenient dependency grounded in some property $p$. Then $p$ must be continually present during the detection of consciousness: that is, there can be no temporal gaps during which a system $s$ lacks $p$ while $T$ makes non-trivial predictions that $s$ is conscious.
@@
Step 2. Suppose there exists a gap: a time $t^{*}$ at which $s$ lacks $p$ but $T$ still predicts that $s$ is conscious.
Step 3. At $t^{*}$, without $p$, there is no lenient dependency between predictions and inferences for $T$ regarding $s$.
Step 4. If there is no lenient dependency, both horns of the Kleiner-Hoel dilemma apply. Either (a) $T$ grounds predictions in properties that a universal substitution can negate or change significantly, all while inferences remain fixed, creating mismatches, in which case $T$ is a priori falsified; or (b) $T$ grounds predictions in input/output structure alone, creating strict dependency with inferences.
Step 5. If (a) the theory is falsified, and if (b), then by Definition 4.1 of Trivial Theories, $T$ is trivial for predictions at $t^{*}$. ∎
The only exception to this would be if both (a) a theory could be tested only under conditions of lenient dependency and those instances are enough scientific weight to carry a theory, and also (b) the theory were grounded in different properties at different times. However, this seems to indicate a handful of separable theories of consciousness, and the theory that was not-grounded in continual lenient dependency would fail to pass through the Kleiner-Hoel dilemma.
@@
That is, for any case of lenient dependency, it must be constantly ongoing during predictions of consciousness. This significantly constrains theories of consciousness. It also makes it very clear that it is not the kind of system being tested that matters (e.g., if a system is plastic, or reentrant, etc.) but rather necessarily the kind of properties a theory of consciousness is itself in grounded in all the time.
@@ 5.1 Continual Learning as Continual Lenient Dependency
Lenient dependency requires that a theory’s predictions not be a priori falsified by universal substitutions but it is also not strictly dependent, and we now know from Proposition 5.1 that this must continually hold or else a theory of consciousness collapses. Here, I show that focusing on continual learning offers such a hypothesis.
@@
    Contain internal memory: the static universal substitute tracks history internally to match the learning system’s input/output function. But a system with internal states that update based on inputs, and so change the overall input/output function, is itself a learning system. Therefore, this does not constitute a non-learning substitution.
@@
    Contain external memory: requiring input $(x,\text{history})$ rather than $x$ alone, like a large (or even infinite) static lookup table. However, this means that the input/output is not preserved in the universal substitution, since the input is substantially different. It is not, e.g., the same input information but in a different temporal sequence; in cases of input $(x,\text{history})$ there is substantially more information in the substituted system’s input, thus making it a false universal substitution since the input domain has substantial differences (essentially, it is injected with the memory each time).
@@
    Are derivative: The original Substitution Argument conceptualized a substitution as being swapped into an experimental analysis (e.g., a FNN for an RNN). One might ask if a learning system could be swapped for a series of static systems instead, each of which is like a static snapshot of the system. This series might functionally mimic learning. However, testing in such a system cannot be a swap, but rather, an ongoing reconstruction. This means that any such series is continually dependent and cannot be tested in isolation, i.e., it cannot be truly substituted in for the original.
@@
Already, [o2025caveat] has shown a specific case of Proposition 5.3, that plastic RNNs cannot be unfolded into an equivalent static FNN: this includes any infinite series. However, this argument is not a repetition of their argument: by focusing on learning rather than on plasticity, it offers a broader case. E.g., in [o2025caveat] many of the proofs are based on resource constraints, and do not rule out cases like an infinite static system that contains all possible histories acting as a viable substitution (which are physically impossible, but still potentially relevant to reasoning about falsification). However, Proposition 5.3 does, since it shows that input $(x,\text{history})$ violates the substitution. Note that this applies to, e.g., LLMs acting as substitutes for humans: for an LLM, the context window is fed back into the input, thus, LLMs contain external memory (Proposition 5.3.b) and their inputs are actually quite unlike humans even if they were to act as attempted substitutes in, e.g., a Turing Test [turing2021computing] (each query actually involves the entire previous conversation fed into a static system).
@@
Given Proposition 5.3’s truth, the failure of all static systems to offer viable universal substitutions for learning systems means that the set of universal substitutions must be a smaller subset for theories of consciousness that involve learning. In other words, theories of consciousness that involve learning are harder to a priori falsify by finding universal substitutions. Consider how much more it requires than matching static input/output functions: A universal substitution for a learning system needs to somehow match the entire history of the learning process, while still at the same time ensuring equivalency across input/output! This is not to say that these results currently show that all universal substitutions for learning systems are impossible; the clearest option seems to be, e.g., simulations of a system on a computer, which might, depending on the theory, create pathological mismatches. Addressing this is a direction for future research. However, they do rule out an entire class of substitutions; and, most importantly, the class of substitutions focused solely on fixing inputs/outputs alone.
@@
Finally, we can ask whether (and why) theories of consciousness focused on the process of continual learning would not be strictly dependent. This is much easier to show, since it simply has to be the case that pred and inf can vary. This seems the case due to things like the multiple-realizability of learning, or cases of latent learning, and so on.
@@
Step 1. Consider a continual-learning-based theory of consciousness, wherein pred (the prediction function) is grounded in physical plasticity states, while inf (the inference function) is derived from behavioral outputs, including observable evidence of learning.
Step 2. Strict dependence requires the existence of a function $f$ such that $\mathrm{pred}(o)=f(\mathrm{inf}(o))$ for all observations $o\in\bar{O}$.
Step 3. Consider latent learning: a system can undergo plastic changes that do not immediately manifest in behavior. Let $o$ and $o^{\prime}$ be observations from two systems (or the same system at different times) with identical behavioral outputs but different plasticity states.
Step 4. Since behavioral outputs are identical, $\mathrm{inf}(o)=\mathrm{inf}(o^{\prime})$. Since plasticity states differ, $\mathrm{pred}(o)\neq\mathrm{pred}(o^{\prime})$.
Step 5. Therefore, some learning-based theories are not strictly dependent. ∎
Additionally, it is interesting that there is a relationship to classic philosophical issues of phenomenological consciousness vs. access consciousness. Learning is about counterfactuals, generalization, and dispositions, which all involve information far beyond any given particular test. In other words, a theory’s predictions grounded in learning (say, focused on physical plasticity) “overflow” the access of individual experimental inferences, since the subject of those predictions (like plastic changes) implicitly have effects on a much larger domain (the entire set of all possible experiments, not just the current the experiment).
Proposition 5.5 (Learning Overflows Finite Access).
A theory of consciousness grounded in learning overflows the finite access of individual experiments.
Proof.
Let $\mathcal{P}$ be the space of possible input/output functions while $\mathcal{A}\subset\mathcal{P}$ are the actually tested input/output, focused on a particular experience, given a theory of consciousness.
Step 1. Learning modifies the system’s dispositional structure—its response is across $\mathcal{P}$. Thus $\mathrm{pred}$, while it may be focused on the process of ongoing plasticity or change underlying learning during an individual experiment, has longterm effects over $\mathcal{P}$.
Step 2. Standard empirical inference about consciousness accesses observed behaviors only across the smaller domain of $\mathcal{A}$. That is, traditionally, $\mathrm{inf}$ is defined over $\mathcal{A}$ (just looking at reports or behavior or output in immediate response to stimuli).
Step 3. Since $\mathcal{A}\subset\mathcal{P}$ (proper subset), the domains differ.
Step 4. Functions with different domains cannot be identical. Therefore $\mathrm{pred}\neq\mathrm{inf}$. ∎
It is worth noting that this looks close to the philosophical debate around whether phenomenological consciousness “overflows” access consciousness [block2011perceptual].

### v2

@@ 4 The Proximity Argument
Zooming out: there exist in the literature previous “no-go” results on AI consciousness, such as arguments against AI consciousness based on causal exclusion [46] (though similar arguments have been given against the efficacy of human consciousness as well [44]), as well as other arguments for biological naturalism [66, 21]. However, one unique aspect of the Proximity Argument is that it makes minimal assumptions other than that: (a) a theory of consciousness cannot be trivial or a priori falsified, (b) that certain universal substitutions are possible (which we know from existent mathematical theorems to be true), and (c) that of widespread or universal mismatches from a theory in the comparative testing setup described would falsify that theory. Nor does it rely on a particular theory; e.g., while transformers provably are entirely feedforward and have zero integrated information (thus, IIT would say that LLMs are not conscious) [5], the Proximity Argument goes beyond relying on a particular theory. It requires no opinion on whether qualia is causally relevant or epiphenomenal, nor whether functionalism is true, nor any other philosophical position about consciousness, nor does it rely on any particular theory of consciousness or even class of theories to be true (it does not, e.g., require solving the Hard Problem [22]). It is a philosophically neutral proof that both constrains and disproves LLM consciousness based on falsification, linking up to historical critiques about functionalism and intelligence (e.g., the Blockhead thought experiment) [16].
It should also be noted that, following Theorem 3.1, aspects of the Proximity Argument could be restated in terms of the shortest program $K(f)$ acting as the substitute for the LLM. Indeed, the shortest program is much like a mirror of a lookup table: instead of being the maximally-complex implementation (wherein every part relates to the function, at least), the shortest program is the minimally-complex implementation, and it too is (a) different from whatever original program or system was specified, and (b) just as much a direct function of input/output.
4.1 The Proximity Argument is Particularly Forceful for LLMs
@@
The Proximity Argument is based on the formal falsification framework developed in [45]; however, while applying it to LLMs, the Proximity Argument does not necessarily require the full acceptance of the Kleiner-Hoel dilemma applying in humans for all possible theories of consciousness. Humans are much less proximate (Proposition 4.5) in “substitution distance” (Definition 4.3) to trivially non-conscious systems, therefore the Constraint Theorem (Theorem 4.6) doesn’t bound the possibility space to the same degree, as there are many more properties lost in universal substitutions. This means there are far more available choices to ground non-trivial theories of consciousness in (although whether that can escape both horns of the dilemma is another matter). At least, with humans, we have good self-introspective reasons to believe we are conscious; theories grounded in phenomenology may even be different from normal scientific theories due to us having first-person evidence [67]. However, grounding theories in phenomenology alone likely requires sacrifices (like falsifiability), and, in light of the following evidence that some theories of consciousness do escape logical issues around falsifiability, this option is set aside here.
A related objection is that we should artificially restrict our scope of consciousness science to apply solely in humans to avoid the Kleiner-Hoel dilemma [3, 72]. This too comes with significant costs, like: (a) substitutions may be harder to a priori identify, but will still apply in humans simply due to independence between inferences and predictions (e.g., [29]), since the difference between a “substitution” and an “advanced experiment” seems impossible to specify; (b) theories lose their universality and this calls into question any extrapolation beyond humans, which represents a large domain for which a science of consciousness should have answers; (c) it is equivalent to saying that, beyond humans, inferences based on general conscious-like behavior contain zero actionable scientific information and cannot be used for inferences about consciousness.
It is (c) that is actually most relevant here, as it makes the Proximity Argument particularly strong: If one tried to get around the Proximity Argument by arguing to solely confirm/falsify theories of consciousness in humans to try to preempt issues of falsification (all while ignoring (a), that substitutions still likely apply), this move would necessarily imply that the intelligent outputs of LLMs contain zero information about their consciousness to begin with and cannot be used for inferences; therefore, we would have no reason to believe LLMs are conscious.
Luckily, these debates can be sidestepped entirely by examples of lenient dependency that avoid the Kleiner-Hoel dilemma in humans but not LLMs. If this can be shown to hold, meaning that we have reasons to believe the Kleiner-Hoel dilemma doesn’t apply to humans under particular classes of theories of consciousness, we have even stronger reasons to think the Proximity Argument rules out consciousness in LLMs, since there can be no argument that it rules out consciousness research altogether.
Therefore, the second part of this paper will explore the requirements for theories to avoid the Kleiner-Hoel dilemma in humans, a property herein referred to as “lenient dependency.”
Definition 4.11 (Lenient Dependency).
A theory of consciousness $T$ satisfies lenient dependency if there is not strict dependency between its predictions and inferences and it has no pathological mismatches in predictions due to definable possible universal substitutions.
@@
First, I argue that lenient dependency must be a continually-existing criterion for theories of consciousness (a high bar). Then, I propose that learning is a form of lenient dependency in humans, and that this is necessarily restricted to truly continual learning (the learning occurring with, or potentially even constituting, every single conscious experience). If this further argument is correct, it would significantly constrain consciousness in humans to involve the process of continual learning. I argue that this restriction is scientifically positive: while it rules out most existing theories in the field, it still points us toward evocative and rich theories of consciousness and reveals what a minimal falsifiable and non-trivial theory of consciousness would look like. Additionally, it indicates that a fundamental difference between LLMs and humans is the ability to continually learn and that LLM’s lack of consciousness may be tied to their static nature.
5 The Necessity of Continual Lenient Dependency
@@
Here, I argue that avoiding the Kleiner-Hoel dilemma forms a strong constraint on theories, in that theories must continually be based on properties that satisfy lenient dependency at every prediction by the theory with regard to the contents or presence of consciousness. In other words, lenient dependency is a continual and ongoing requirement for theories of consciousness.
@@
Let $T$ be a theory of consciousness with lenient dependency grounded in some property $p$. Then $p$ must be continually present during the detection of consciousness: that is, there can be no temporal gaps during which a system $s$ lacks $p$ while $T$ continues to make non-trivial predictions that $s$ is conscious.
@@
Step 2. Suppose there exists a gap: a time $t^{*}$ at which $s$ lacks $p$, yet $T$ still predicts that $s$ is conscious.
Step 3. At $t^{*}$, without $p$, system $s$ can be substituted with a non-conscious system $s^{\prime}$ (Definition 4.2), such as a static single-hidden-layer FNN (Corollary 4.10).
Step 4. Therefore, $T$ must either (a) make different predictions for $s^{\prime}$, creating pathological mismatches, or (b) predict consciousness in $s^{\prime}$, indicating that $T$ itself is trivial.
The only exception to this would be if both (a) a theory could be tested only under conditions of lenient dependency and those instances are enough scientific weight to carry a theory, and also (b) the theory were grounded in different properties at different times. However, this seems to indicate two separable theories of consciousness, and the theory that was not grounded in continual lenient dependency would fail for the same reasons as in Proposition 5.1.
@@
That is, for any case of lenient dependency, it must be constantly ongoing during predictions of consciousness. This significantly constrains theories of consciousness. It also makes it very clear that it is not the kind of system being tested that matters (e.g., if a system is plastic, or reentrant, etc.) but rather the properties a theory of consciousness is itself grounded in all the time.
@@ 5.1 Continual Learning as Continual Lenient Dependency
Lenient dependency requires that a theory’s predictions not be a priori falsified by known universal substitutions but is also not strictly dependent, and we now know from Proposition 5.1 that this must continually hold or else a theory of consciousness collapses. Here, I show that a theory involving continual learning offers such a hypothesis.
@@
    Contain internal memory: the static universal substitute may track history internally to match any given learning system’s input/output function. But by definition, any system with internal states that itself updates based on inputs, and so change its overall input/output function, is also itself a learning system. Therefore, this does not constitute a non-learning substitution.
@@
    Contain external memory: requiring input $(x,\text{history})$ rather than $x$ alone, like a large (or even infinite) static lookup table. However, this requires that the input/output is not preserved in the universal substitution, since the input is substantially different. It is not, e.g., the same input information but in a different temporal sequence; in cases of input $(x,\text{history})$ there can be orders of magnitude greater information in the substituted system’s input, thus making it a false universal substitution over the input domain.
@@
    Be derivative: One might ask if a learning system could be substituted for a series of static systems instead, each of which is like a static snapshot of the system. This series might functionally mimic learning. However, testing in such a system cannot be a swap, but rather, an ongoing reconstruction. This means that any such series cannot be tested in isolation, i.e., it cannot be truly substituted in for the original, as it depends on monitoring changes in the original.
@@
Already, [59] has shown a specific case of Proposition 5.3, that plastic RNNs cannot be unfolded into an equivalent static FNN. However, this argument is not a repetition of their argument: by focusing on learning rather than on plasticity, and substitutions in general rather than just unfolded networks, it offers a broader case. E.g., in [59] many of the proofs are based on resource constraints, and do not rule out cases like an infinite static system that contain all possible histories acting as a viable substitution (which are physically impossible, but still potentially relevant to reasoning about falsification). However, Proposition 5.3 does, since it shows that input $(x,\text{history})$ violates the substitution. Note that this applies to, e.g., LLMs acting as substitutes for humans: for an LLM, the context window is fed back into the input, thus, LLMs contain external memory (Proposition 5.3(b)) and their inputs are actually quite unlike humans even if they were to act as substitutes in, e.g., a Turing Test [71] (each query actually involves the entire previous conversation fed into a static system).
@@
Given Proposition 5.3’s truth, the failure of all static systems to offer viable universal substitutions for learning systems means that the set of universal substitutions is likely a smaller subset when it comes to theories of consciousness that involve learning. In other words, theories of consciousness that involve learning are harder to a priori falsify by finding universal substitutions. Consider how much more it requires than matching static input/output functions: A universal substitution for a learning system needs to somehow match the entire history of the learning process, while still at the same time ensuring equivalency across input/output! This is not to say that these results currently show that all universal substitutions for learning systems are impossible; the clearest option seems to be, e.g., simulations of a system on a computer, which might, depending on the theory, create pathological mismatches. Addressing this is a direction for future research. However, they do rule out the entire class of static substitutions for theories of consciousness within which predictions are grounded in or require continual learning.
@@
Finally, we can ask whether (and why) theories of consciousness focused on the process of continual learning would not be strictly dependent. This is much easier to show, since to hold it simply has to be the case that pred and inf can vary or don’t stem from the exact same data. This seems likely the case due to things like the multiple-realizability of learning, or cases of latent learning, and so on.
@@
Step 1. Consider a continual-learning-based theory of consciousness, wherein pred is grounded in the ongoing process of plasticity, while inf is derived from behavioral outputs that evince observable evidence of learning.
Step 2. Consider latent learning: a system can undergo plastic changes that do not immediately manifest in behavior. Let $o$ and $o^{\prime}$ be observations from two systems with identical behavioral outputs (a static substitution) but different plasticity states.
Step 3. Since behavioral outputs are identical, $\mathrm{\textit{inf}}(o)=\mathrm{\textit{inf}}(o^{\prime})$. However, since plasticity states differ, $\mathrm{\textit{pred}}(o)\neq\mathrm{\textit{pred}}(o^{\prime})$.
Step 4. If $\mathrm{\textit{pred}}(o)\neq\mathrm{\textit{pred}}(o^{\prime})$ following a static substitution then strict dependency cannot hold for the theory. ∎
Finally, it is worth noting that there exists an obvious relationship regarding the overflow of predictions with regards to inferences to the classic philosophical issues of phenomenological consciousness vs. access consciousness [17]. Learning is about counterfactuals, generalization, and dispositions, which all involve information far beyond any given particular test. In other words, a theory’s predictions grounded in learning (say, focused on physical plasticity) “overflow” the access of individual experimental inferences, since the subject of those predictions (like plastic changes) implicitly have effects on a much larger domain (the entire set of all possible experiments, not just the current experiment).